OpenStudy (anonymous):

Solve 5^x= 81.2 Round answer to the nearest ten-thousandth x=_______

OpenStudy (solomonzelman):

\[ 5^x= 81.2\]\[ \log~5^x= \log~81.2\]\[x~ \log~5= \log~81.2\]\[x~=\frac{ \log~81.2}{\log5}\]can you do it from here ?

OpenStudy (anonymous):

answer would be 2.731?

OpenStudy (anonymous):

is that to the ten thousandth place?

OpenStudy (anonymous):

x≈2.7320

OpenStudy (anonymous):

thank you babe

OpenStudy (solomonzelman):

I got to go, have a good night. You were Atlanta GA just now...

OpenStudy (anonymous):

cool

OpenStudy (solomonzelman):

;)

OpenStudy (anonymous):

<3